# Arithmetical enhancements of the Kogbetliantz method for the SVD of order two

### v1

@@ 2 Floating-point considerations
For the second equality in (2), note that $(\nu+1)^{2}=\nu^{2}+1+2\nu\geq\nu^{2}+1\geq\nu^{2}$ since $\nu>0$. By taking the square roots, it follows that $\nu+1\geq\sqrt{\nu^{2}+1}\geq\nu$, and therefore
@@
An underlined expression denotes a computed floating-point approximation of the exact value of that expression. Given $\tan(2\phi)$, for $40\leq|\phi|\leq\pi/$ 4, $\tan\phi$ and $\underline{\tan\phi}$ are
$\tan\phi=\frac{\tan(2\phi)}{1+\sqrt{\tan^{2}(2\phi)+1}},\qquad\underline{\tan\phi}=\mathop{\mathrm{fl}}\left(\frac{\mathop{\mathrm{fl}}(\tan(2\phi))}{\mathop{\mathrm{fl}}(1+\mathop{\mathrm{hypot}}(\mathop{\mathrm{fl}}(\tan(2\phi)),1))}\right),$ (3)
if $\tan(2\phi)$ and $\mathop{\mathrm{fl}}(\tan(2\phi))$ are finite, or $\mathop{\mathrm{sign}}(\tan(2\phi))$ and $\mathop{\mathrm{sign}}(\mathop{\mathrm{fl}}(\tan(2\phi)))$ otherwise.
@@
Let $\tan(2\phi)$ be given, such that $\mathop{\mathrm{fl}}(\tan(2\phi))$ is finite. Then, under the assumptions of Lemma 2.1, for $\underline{\tan\phi}$ from (3) holds $0\leq|\underline{\tan\phi}|\leq 1$.
@@
For $|\mathop{\mathrm{fl}}(\tan(2\phi))|=\nu$, due to Lemma 2.1, $\mathop{\mathrm{hypot}}(\mathop{\mathrm{fl}}(\tan(2\phi)),1)=\nu$, and so the denominator in (3) is $\mathop{\mathrm{fl}}(1+\nu)=\nu$. Note that the numerator is always at most as large in magnitude as the denominator. Thus, $0\leq|\underline{\tan\phi}|\leq 1$, what had to be proven. ∎
@@
Let $\tan{\phi}$ be given, for $|\phi|\leq\pi/2$. Then, $\sec{\phi}=1/\cos{\phi}$ can be approximated as $\underline{\sec{\phi}}=\mathop{\mathrm{hypot}}(\mathop{\mathrm{fl}}(\tan{\phi}),1)$. If $\tan{\phi}=\mathop{\mathrm{fl}}(\tan{\phi})$, then $\underline{\sec{\phi}}=\mathop{\mathrm{fl}}(\sec{\phi})$. When the assumptions of Lemma 2.1 hold and $\mathop{\mathrm{fl}}(\tan{\phi})$ is finite, so is $\underline{\sec{\phi}}$.
@@
The approximation relation follows from the definition of $\mathop{\mathrm{hypot}}$ and from $\sec\phi=\sqrt{\tan^{2}\phi+1}$, while its finiteness for a finite $\mathop{\mathrm{fl}}(\tan\phi)$ follows from Lemma 2.1, since $|\mathop{\mathrm{fl}}(\tan\phi)|\leq\nu$ implies $\mathop{\mathrm{hypot}}(\mathop{\mathrm{fl}}(\tan\phi),1)\leq\nu$. ∎
@@
In the second case in (5), $c>0$ and $\mu\leq|z|\leq\nu$, if $e_{\nu}\geq e_{\mu}+2p-1$.
@@ 3 The SVD of general matrices of order two
For $\mathfrak{s}=0$, there is one equivalence class of matrix types, represented by $\mathfrak{t}=9$. For $\mathfrak{s}=1$, there are three classes, represented by $\mathfrak{t}=3$, $\mathfrak{t}=5$, and $\mathfrak{t}=13$, while for $\mathfrak{s}=2$ there is one class, $\mathfrak{t}=15$. The SVD computation for the first three classes is straightforward, while for the fourth and the fifth class is more involved. A matrix of any type, except $\mathfrak{t}=15$, can be permuted into an upper triangular one. If a matrix so obtained is well scaled, its SVD can alternatively be computed by xLASV2. However, xLASV2 does not accept general matrices (i.e., $\mathfrak{t}=15$), unlike the proposed method, which is a modification of [15] when $J=I$, and consists of the following three phases:
@@ 3.1 Prescaling of the matrix and the simple cases ($\mathfrak{t}\cong 3,5,9$)
$\epsilon_{\theta}^{\prime}=\frac{\tan^{2}\theta}{\tan^{2}\theta+1}(\delta_{\theta}^{2}-1),\qquad 0<\frac{\tan^{2}\theta}{\tan^{2}\theta+1}\leq\frac{1}{2}.$
@@ 3.2.2 A fully pivoted URV when $\mathfrak{t}^{\prime}=15$
The same holds if $x$ and $y$ are of the opposite signs and the subtraction is taken instead. Specifically, from (20), the bound (26) holds for $x^{\prime}=\underline{g_{12}^{\prime\prime\prime}}$ and $y^{\prime}=\underline{g_{22}^{\prime\prime\prime}}$ of the same sign, with $\pm^{\prime}=+$, and for $x^{\prime\prime}=\underline{g_{22}^{\prime\prime\prime}}$ and $y^{\prime\prime}=\underline{g_{12}^{\prime\prime\prime}}$ of the opposite signs, with $\pm^{\prime\prime}=-$.
@@
$\mathop{\mathrm{fl}}\left(\frac{\mathop{\mathrm{fma}}(\pm^{!}y^{!},\underline{\tan\vartheta},x^{!})}{\underline{\sec\vartheta}}\right)=\frac{x^{!}\pm^{!}y^{!}\tan\vartheta}{\sec\vartheta}\cdot\frac{\delta_{\mathop{\mathrm{fma}}}\delta_{/}^{\prime\prime}}{\delta_{\vartheta}^{\prime}}\cdot\delta_{\pm}=\underline{r_{?2}^{\prime\prime}}\cdot\delta_{1}^{!}\cdot\delta_{\pm}=\underline{r_{?2}^{\prime\prime}}\delta_{2}^{!},$ (28)
where $!=^{\prime}$ for ?$=1$ and $!=^{\prime\prime}$ for ?$=2$. Now (23) follows from bounding $\delta_{1}^{!}=\delta_{\mathop{\mathrm{fma}}}\delta_{/}^{\prime\prime}/\delta_{\vartheta}^{\prime}$ below and above using (16). Since $\delta_{2}^{!}=\delta_{1}^{!}\delta_{\pm}$ in (28), (24) is a consequence of (23). ∎
@@
$\frac{(1-\varepsilon_{\text{{T}}})^{2}(1-\varepsilon)^{2}}{\sqrt{((1+\varepsilon)^{2}+1)/2}(1+\varepsilon)}=\delta_{3}^{-}\leq\delta_{3}^{\prime},\delta_{3}^{\prime\prime}\leq\delta_{3}^{+}=\frac{(1+\varepsilon_{\text{{T}}})^{2}(1+\varepsilon)^{2}}{\sqrt{((1-\varepsilon)^{2}+1)/2}(1-\varepsilon)}.$ (31)
@@
+¯$\underline{\star_{+}}&=\mathop{\mathrm{FL}}(\star_{+})=\star_{+}(1+\epsilon_{+}^{\prime}),\quad\star_{+}=g_{12}^{\prime\prime\prime}g_{11}^{\prime\prime\prime}+g_{22}^{\prime\prime\prime}g_{21}^{\prime\prime\prime},\quad|\epsilon_{+}^{\prime}|\leq\varepsilon_{\text{{T}}},\\ \underline{\star_{-}}&=\mathop{\mathrm{FL}}(\star_{-})=\star_{-}(1+\epsilon_{-}^{\prime}),\quad\star_{-}=g_{22}^{\prime\prime\prime}g_{11}^{\prime\prime\prime}-g_{12}^{\prime\prime\prime}g_{21}^{\prime\prime\prime},\quad|\epsilon_{-}^{\prime}|\leq\varepsilon_{\text{{T}}},\end{aligned}\\ \mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime})=(\star_{\pm}/g_{11}^{\prime\prime\prime})(1+\epsilon_{\pm}^{\prime})(1+\epsilon_{/}^{\prime\prime\prime}),\quad|\epsilon_{/}^{\prime\prime\prime}|\leq\varepsilon_{\text{{T}}}.\end{gathered}$ (32)
The quotient $\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime})$, converted to the working precision with $\mathop{\mathrm{fl}}(\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}))$, is possibly not correctly rounded from the value of ±¯$\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}$ due to its double rounding. Using the assumption that no underflow in the working precision occurs, it follows that
$\mathop{\mathrm{fl}}(\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}))=(\star_{\pm}/g_{11}^{\prime\prime\prime})(1+\epsilon_{\pm}^{\prime})(1+\epsilon_{/}^{\prime\prime\prime})(1+\epsilon_{\mathop{\mathrm{fl}}})=(\star_{\pm}/g_{11}^{\prime\prime\prime})\delta_{\pm}^{\prime},\quad|\epsilon_{\mathop{\mathrm{fl}}}|\leq\varepsilon,$ (33)
with $\delta_{\pm}^{\prime}=(1+\epsilon_{\pm}^{\prime})(1+\epsilon_{/}^{\prime\prime\prime})(1+\epsilon_{\mathop{\mathrm{fl}}})$. For the final division by $\underline{\sec\vartheta}$, due to (16), holds
@@
By fixing the sign in the $\pm$ subscript in (32), (33), and (34), let $\delta_{3}^{\prime}=\delta_{-}^{\prime}\delta_{/}^{\prime\prime\prime\prime}/\delta_{\vartheta}^{\prime}$ and $\delta_{3}^{\prime\prime}=\delta_{+}^{\prime}\delta_{/}^{\prime\prime\prime\prime}/\delta_{\vartheta}^{\prime}$. Now bound $\delta_{3}^{\prime}$ and $\delta_{3}^{\prime\prime}$ below by $\delta_{3}^{-}$ and above by $\delta_{3}^{+}$, by combining the appropriate lower and upper bounds for $\delta_{\pm}^{\prime}$, $\delta_{\vartheta}^{\prime}$ from Lemma 3.1, and $\delta_{/}^{\prime\prime\prime\prime}$, where
$\frac{(1-\varepsilon_{\text{{T}}})^{2}(1-\varepsilon)\cdot(1-\varepsilon)}{\sqrt{((1+\varepsilon)^{2}+1)/2}(1+\varepsilon)}=\delta_{3}^{-}\leq\delta_{3}^{\prime},\delta_{3}^{\prime\prime}\leq\delta_{3}^{+}=\frac{(1+\varepsilon_{\text{{T}}})^{2}(1+\varepsilon)\cdot(1+\varepsilon)}{\sqrt{((1-\varepsilon)^{2}+1)/2}(1-\varepsilon)},$
@@ 4 Numerical testing
A timing comparison of xLASV2 and the proposed method was not performed since the correctly rounded routines are still in development. By construction the proposed method is more computationally complex than xLASV2, so it is expected to be slower.

### v2

@@ 2 Floating-point considerations
For the second equality in (2), note that $(\nu+1)^{2}=\nu^{2}+1+2>\nu^{2}+1>\nu^{2}$ since $\nu>0$. By taking the square roots, it follows that $\nu+1>\sqrt{\nu^{2}+1}>\nu$, and therefore
@@
An underlined expression denotes a computed floating-point approximation of the exact value of that expression. Assuming $40\leq|\phi|\leq\pi/$ 4, and with $\tan(2\phi)$ and $\underline{\tan(2\phi)}$ given, where the latter may or may not be equal to $\mathop{\mathrm{fl}}(\tan(2\phi))$, $\tan\phi$ and $\underline{\tan\phi}$ are
$\tan\phi=\frac{\tan(2\phi)}{1+\sqrt{\tan^{2}(2\phi)+1}},\qquad\underline{\tan\phi}=\mathop{\mathrm{fl}}\left(\frac{\underline{\tan(2\phi)}}{\mathop{\mathrm{fl}}(1+\mathop{\mathrm{hypot}}(\underline{\tan(2\phi)},1))}\right),$ (3)
if $\tan(2\phi)$ and $\underline{\tan(2\phi)}$ are finite, or $\mathop{\mathrm{sign}}(\tan(2\phi))$ and $\mathop{\mathrm{sign}}(\underline{\tan(2\phi)})$ otherwise.
@@
Let $\tan(2\phi)$ be given, such that $\underline{\tan(2\phi)}$ is finite. Then, under the assumptions of Lemma 2.1, for $\underline{\tan\phi}$ from (3) holds $0\leq|\underline{\tan\phi}|\leq 1$.
@@
For $|\underline{\tan(2\phi)}|=\nu$, due to Lemma 2.1, $\mathop{\mathrm{hypot}}(\underline{\tan(2\phi)},1)=\nu$, and so the denominator in (3) is $\mathop{\mathrm{fl}}(1+\nu)=\nu$. Note that the numerator is always at most as large in magnitude as the denominator. Thus, $0\leq|\underline{\tan\phi}|\leq 1$, what had to be proven. ∎
@@
Let $\tan{\phi}$ be given, for $|\phi|\leq\pi/2$. Then, $\sec{\phi}=1/\cos{\phi}$ can be approximated as $\underline{\sec{\phi}}=\mathop{\mathrm{hypot}}(\underline{\tan{\phi}},1)$. If $\tan{\phi}=\underline{\tan{\phi}}$, then $\underline{\sec{\phi}}=\mathop{\mathrm{fl}}(\sec{\phi})$. When the assumptions of Lemma 2.1 hold and $\underline{\tan{\phi}}$ is finite, so is $\underline{\sec{\phi}}$.
@@
The approximation relation follows from the definition of $\mathop{\mathrm{hypot}}$ and from $\sec\phi=\sqrt{\tan^{2}\phi+1}$, while its finiteness for a finite $\underline{\tan\phi}$ follows from Lemma 2.1, since $|\underline{\tan\phi}|\leq\nu$ implies $\mathop{\mathrm{hypot}}(\underline{\tan\phi},1)\leq\nu$. ∎
@@
If $e_{\nu}\geq e_{\mu}+2p-1$, there exists $c>0$ such that $\mu\leq|z|\leq\nu$ in (5).
@@ 3 The SVD of general matrices of order two
For $\mathfrak{s}=0$, there is one equivalence class of matrix types, represented by $\mathfrak{t}=9$. For $\mathfrak{s}=1$, there are three classes, represented by $\mathfrak{t}=3$, $\mathfrak{t}=5$, and $\mathfrak{t}=13$, while for $\mathfrak{s}=2$ there is one class, $\mathfrak{t}=15$. The SVD computation for the first three classes is straightforward, while for the fourth and the fifth class is more involved. A matrix of any type, except $\mathfrak{t}=15$, can be permuted into an upper triangular one. If a matrix so obtained is well scaled, its SVD can alternatively be computed by xLASV2. However, xLASV2 does not accept general matrices (i.e., $\mathfrak{t}=15$), unlike the proposed method, which is a modification of the “trigonometric” case from [15], i.e., of the case where the sign matrix $J=I$. The proposed method consists of the following three phases:
@@ 3.1 Prescaling of the matrix and the simple cases ($\mathfrak{t}\cong 3,5,9$)
$\epsilon_{\theta}^{\prime}=\frac{\tan^{2}\theta}{\tan^{2}\theta+1}(\delta_{\theta}^{2}-1),\qquad 0<\theta\leq\frac{\pi}{4}\implies 0<\frac{\tan^{2}\theta}{\tan^{2}\theta+1}\leq\frac{1}{2}.$
@@ 3.2.2 A fully pivoted URV when $\mathfrak{t}^{\prime}=15$
The same holds if $x$ and $y$ are of the opposite signs and the subtraction is taken instead. Namely, from (20), the bound (26) holds for $x^{\prime}=\underline{g_{12}^{\prime\prime\prime}}$ and $y^{\prime}=\underline{g_{22}^{\prime\prime\prime}}$ of the same sign,
$\mathop{\mathrm{sign}}(x^{\prime}=\underline{g_{12}^{\prime\prime\prime}})=\mathop{\mathrm{sign}}(y^{\prime}=\underline{g_{22}^{\prime\prime\prime}})\implies\delta_{\pm}\mapsto\delta_{+},$ (27)
i.e., with $\delta_{\pm}$ denoted as $\delta_{+}$, and for $x^{\prime\prime}=\underline{g_{22}^{\prime\prime\prime}}$ and $y^{\prime\prime}=\underline{g_{12}^{\prime\prime\prime}}$ of the opposite signs,
$\mathop{\mathrm{sign}}(x^{\prime\prime}=\underline{g_{22}^{\prime\prime\prime}})=-\mathop{\mathrm{sign}}(y^{\prime\prime}=\underline{g_{12}^{\prime\prime\prime}})\implies\delta_{\pm}\mapsto\delta_{-},$ (28)
with $\delta_{\pm}$ becoming $\delta_{-}$ in this case. When both cases are considered together, $\delta_{\pm}$ is used.
@@
$\mathop{\mathrm{fl}}\left(\frac{\mathop{\mathrm{fma}}(y^{\prime},\underline{\tan\vartheta},x^{\prime})}{\underline{\sec\vartheta}}\right)=\frac{x^{\prime}+y^{\prime}\tan\vartheta}{\sec\vartheta}\cdot\frac{\delta_{\mathop{\mathrm{fma}}}\delta_{/}^{\prime\prime}}{\delta_{\vartheta}^{\prime}}\cdot\delta_{+}=\underline{r_{12}^{\prime\prime}}\cdot\delta_{1}^{\prime}\cdot\delta_{+}=\underline{r_{12}^{\prime\prime}}\delta_{2}^{\prime},$ (30)
in the case (27), while in the case (28) it is similarly obtained that
$\mathop{\mathrm{fl}}\left(\frac{\mathop{\mathrm{fma}}(-y^{\prime\prime},\underline{\tan\vartheta},x^{\prime\prime})}{\underline{\sec\vartheta}}\right)=\frac{x^{\prime\prime}-y^{\prime\prime}\tan\vartheta}{\sec\vartheta}\cdot\frac{\delta_{\mathop{\mathrm{fma}}}\delta_{/}^{\prime\prime}}{\delta_{\vartheta}^{\prime}}\cdot\delta_{-}=\underline{r_{22}^{\prime\prime}}\cdot\delta_{1}^{\prime\prime}\cdot\delta_{-}=\underline{r_{22}^{\prime\prime}}\delta_{2}^{\prime\prime}.$ (31)
Now (23) follows from bounding $\delta_{1}^{\prime}$ and $\delta_{1}^{\prime\prime}$, standing for $\delta_{\mathop{\mathrm{fma}}}\delta_{/}^{\prime\prime}/\delta_{\vartheta}^{\prime}$ in (30) and (31), respectively, below and above using (16). Since $\delta_{2}^{\prime}=\delta_{1}^{\prime}\delta_{+}$ in (30), and $\delta_{2}^{\prime\prime}=\delta_{1}^{\prime\prime}\delta_{-}$ in (31), the proof is concluded by observing that (24) is a consequence of (23). ∎
@@
$\frac{(1-\varepsilon_{\text{{T}}})(1-\varepsilon)^{2}}{\sqrt{((1+\varepsilon)^{2}+1)/2}(1+\varepsilon)}=\delta_{3}^{-}\leq\delta_{3}^{\prime},\delta_{3}^{\prime\prime}\leq\delta_{3}^{+}=\frac{(1+\varepsilon_{\text{{T}}})(1+\varepsilon)^{2}}{\sqrt{((1-\varepsilon)^{2}+1)/2}(1-\varepsilon)}.$ (34)
@@
+¯$\underline{\star_{+}}&=\mathop{\mathrm{FL}}(\star_{+})=\star_{+}(1+\epsilon_{+}^{\prime}),\quad\star_{+}=g_{12}^{\prime\prime\prime}g_{11}^{\prime\prime\prime}+g_{22}^{\prime\prime\prime}g_{21}^{\prime\prime\prime},\quad|\epsilon_{+}^{\prime}|\leq\varepsilon_{\text{{T}}},\\ \underline{\star_{-}}&=\mathop{\mathrm{FL}}(\star_{-})=\star_{-}(1+\epsilon_{-}^{\prime}),\quad\star_{-}=g_{22}^{\prime\prime\prime}g_{11}^{\prime\prime\prime}-g_{12}^{\prime\prime\prime}g_{21}^{\prime\prime\prime},\quad|\epsilon_{-}^{\prime}|\leq\varepsilon_{\text{{T}}},\end{aligned}\\ \mathop{\mathrm{fl}}(\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}))=\mathop{\mathrm{fl}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime})=(\star_{\pm}/g_{11}^{\prime\prime\prime})(1+\epsilon_{\pm}^{\prime})(1+\epsilon_{/}^{\prime\prime\prime})=(\star_{\pm}/g_{11}^{\prime\prime\prime})\delta_{\pm}^{\prime},\end{gathered}$ (35)
with $|\epsilon_{/}^{\prime\prime\prime}|\leq\varepsilon$, since the doubly-rounded quotient $\mathop{\mathrm{fl}}(\mathop{\mathrm{FL}}(\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}))$ is correctly rounded to the working precision from the value of ±¯$\underline{\star_{\pm}}/g_{11}^{\prime\prime\prime}$, where $\star_{\pm}$ can be replaced by $\star_{+}$ or $\star_{-}$. This double rounding is innocuous with the standard datatypes, i.e, when rounding from exact to double to single and from exact to quadruple to double precision, due to [26, Theorem 4] and [27, Theorem 29 and Table II], but might not be for specialized, non-standard datatypes, for which the proposed method would have to be redesigned.
For the final division by $\underline{\sec\vartheta}$, due to (16) and (35), it holds
@@
By fixing the sign to $+$ or -- in the $\pm$ subscripts in (35) and (36), let $\delta_{3}^{\prime}=\delta_{-}^{\prime}\delta_{/}^{\prime\prime\prime\prime}/\delta_{\vartheta}^{\prime}$ and $\delta_{3}^{\prime\prime}=\delta_{+}^{\prime}\delta_{/}^{\prime\prime\prime\prime}/\delta_{\vartheta}^{\prime}$. Now bound $\delta_{3}^{\prime}$ and $\delta_{3}^{\prime\prime}$ below by $\delta_{3}^{-}$ and above by $\delta_{3}^{+}$, by combining the appropriate lower and upper bounds for $\delta_{\pm}^{\prime}$, $\delta_{\vartheta}^{\prime}$ from Lemma 3.1, and $\delta_{/}^{\prime\prime\prime\prime}$, where
$\frac{(1-\varepsilon_{\text{{T}}})(1-\varepsilon)\cdot(1-\varepsilon)}{\sqrt{((1+\varepsilon)^{2}+1)/2}(1+\varepsilon)}=\delta_{3}^{-}\leq\delta_{3}^{\prime},\delta_{3}^{\prime\prime}\leq\delta_{3}^{+}=\frac{(1+\varepsilon_{\text{{T}}})(1+\varepsilon)\cdot(1+\varepsilon)}{\sqrt{((1-\varepsilon)^{2}+1)/2}(1-\varepsilon)},$
@@ 4 Numerical testing
[…]
Table 2: Average run-times per matrix of the methods $\mathtt{K}$ and $\mathtt{L}$, on $\circ$ and $\bullet$ upper triangular random matrices ($22\cdot 2^{30}$ of each kind), and the slowdown of $\mathtt{K}$ versus $\mathtt{L}$.
[…]
$\mathtt{L}$ $1.476\cdot 10^{-6}$ s $1.668\cdot 10^{-6}$ s $1.658\cdot 10^{-6}$ s $1.653\cdot 10^{-6}$ s
[…]